OpenStudy (anonymous):

Calculate the power delivered by a force F acting on a particle that moves with a velocity v, where F = (4.80 N) i + (3.30 N) k and v = (6.50 m/s) i.

OpenStudy (anonymous):

This should help:\[P \overset{\text{def}}{=} \frac{\Delta W}{\Delta t} = \frac{\vec F \cdot \Delta \vec x }{\Delta t} = \vec F \cdot \frac{\Delta \vec x}{\Delta t} = \boxed{\vec F \cdot \vec v}\]